What is the pressure felt at the bottom of a cylindrical tank if the tank has a base radius of 0.5m and a height of 1.75m and the tank carries a fluid with mass of 400kg?
1
Expert's answer
2020-03-16T13:05:28-0400

Mass of fluid"=400Kg"

Tank has a base radius of "0.5m" and a height of "1.75m" .

Volume of tank"=\pi r^2h=\pi\times 0.5^2\times 1.75=1.37\ m^3"

Density of fluid"=\frac{M}{V}=\frac{400}{1.37}=291.17 \ Kg/m^3"

Pressure at bottom"=P_o+\rho gh" (where "P_o" is atmospheric pressure).

"\implies P=(1.01325\times 10^5+291.17\times 9.8\times 1.75)Pa=1.06318\times 10^5 \ Pa=1.049 \ atm"